OpenStudy (anonymous):

URGENT = Prove/ Verify (1-sin(x))/(1+sin(x)) = (sec(x)-tan(x))^2

OpenStudy (anonymous):

are your struggling? :P

OpenStudy (anonymous):

\[\frac{1 - \sin x}{1 + \sin x } \implies \frac{(1-\sin x)^2}{\cos^2 x}\] then you have \[\frac{1 - 2\sin x + \sin^2 x}{\cos^2 x}\] the rest should be algebra

OpenStudy (anonymous):

can you finish the problem please?

OpenStudy (anonymous):

too much hassle on coding the latex :/ can't you see it yet?

OpenStudy (anonymous):

im trying

OpenStudy (anonymous):

\[\frac 1{\cos^2 x} - \frac {2\sin x}{\cos^2 x} + \frac {\sin^2 x}{\cos^2 x}\] i think you already know that 1/cos is sec and sin/cos is tan so the first and last terms i don't have to explain

OpenStudy (anonymous):

OMG I GOT IT

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (anonymous):

ok. welcome